3.5  If an investment of $115,000 resulted in income of $23,000 per year for 10 years, then what was the Rate of Return (ROR), i%, on the investment?  (b) Re-calculate the  Rate of Return (ROR), i%, using the Excel RATE function and record your cell entry and end result.
Answer #1

ANSWER:

via excel:

In excel when we use the rate function we get i = 15.1%

RATE
15.10%

via factor method:

initial investment = $115,000

yearly income = $23,000

n = 10 years

pw = yearly income(p/a,i,n)

115,000 = 23,000(p/a,i,10)

115,000 / 23,000 = (p/a,i,10)

5 = (p/a,i,10)

solving for i via trial and error we get that i is between 15% and 16% via interpolation and subsequently solving further we get that i is 15.1%